OpenStudy (anonymous):

f(x)=(2x-5)(2x+4)(2x+6)/(x+6) find the Asymptotes

OpenStudy (anonymous):

\[\LARGE y=\lim_{x\to \infty}\frac{(2x-5)(2x+4)(2x+6)}{(x+6)}=\] \[\LARGE \lim_{x\to \infty}\frac{(2x-5)(2x+4)(2x+6)}{(x+6)}=\] this is for Vertical asymptote... do you think you can do it? :)

OpenStudy (anonymous):

im scratching my head a bit but uh yes i think i CAN...

OpenStudy (anonymous):

sorry it's not vertical it's horizontal LOL :(

OpenStudy (anonymous):

lol oh ok

OpenStudy (anonymous):

I'll give you another hint.. \[\LARGE \lim_{x\to \infty}\frac{(\frac {2x}{x}-\frac 5x )(\frac{2x}{x}+\frac 4x )(\frac{2x}{x}+\frac 6x )}{\frac xx+\frac 6x }\] ... go on... :)

OpenStudy (anonymous):

then tell me what you get ...

OpenStudy (anonymous):

\[\ \LARGE \lim_{x\to \infty}\frac{(\frac {2x}{x}-\frac 5x )(\frac{2x}{x}+\frac 4x )(\frac{2x}{x}+\frac 6x )}{\frac xx+\frac 6x }=\] \[\LARGE \lim_{x\to \infty}\frac{(2-\frac 5x )(2+\frac 4x )(2+\frac 6x )}{1+\frac 6x }\] \[\LARGE =\frac{(2-0 )(2+0 )(2+0 )}{1+0}=?\]